Let A = {(x, y, z) E R^3ㅣx - y + z = 0}
(a) Show that A is a vector space
(b) Give the dimensions and basis of A.
1
Expert's answer
2021-01-29T16:24:26-0500

(a) We remind that a vector space is a set closed under finite vector addition and scalar multipication.

Since "A\subset{\mathbb{R}}^3" and "{\mathbb{R}}^3" is a vector space, it is enough to check that "A" is a subspace. Assume that "u=(x_1,y_1,z_1),v=(x_2,y_2,z_2)\in A." It means that "x_1-y_1+z_1=0" and "x_2-y_2+z_2=0". We get: "x_1+x_2-(y_1+y_2)+z_1+z_2=0". Thus, "u+v\in A". In addition, for any "\alpha\in{\mathbb{C}}" we have: "\alpha u\in A" ( it holds, because "\alpha x_1-\alpha y_1+\alpha z_1=0" ).Thus, "A" is a vector space.

(b) We remind that a basis is a maximum number of independen vectors. It is obvious that vectors "u=(1,1,0)" and "v=(0,1,1)" belong to "A" and they are independent. We can present any vector "w" that belongs to "A" in the form: "w=(x,x+z,z)". it is obvious that "xu+zv=w". Thus, the dimension of the basis of "A" is 2 and it consists of vectors "u=(1,1,0)" and "v=(0,1,1)".